In a study, 44% of adults questioned reported that their health was excellent. A researcher wishes to study the health of people living close to a nuclear power plant. Find the probability that when 14 adults are randomly selected, more than 3 are in excellent health.

Round to nearest ten-thousandth.  
Answer #1

Let , X be the number of adults that their health was excellent.

Here , X has binomial distribution with parameter n=14 and p=0.44

The pmf of X is ,

P(X=x)=\binom{n}{x}p^xq^{n-x}; x=0,1,2,......,n and q=1-p

= 0 ; otherwise

Now , the probability distribution table is ,

X \binom{n}{x} p^x q^{n-x} P(X=x)
0 1 1 0.000298 0.0003
1 14 0.44 0.000533 0.0033
2 91 0.1936 0.000951 0.0168
3 364 0.08518 0.001699 0.0527
4 1001 0.03748 0.003033 0.1138
5 2002 0.01649 0.005416 0.1788
6 3003 0.00726 0.009672 0.2108
7 3432 0.00319 0.017271 0.1892
8 3003 0.0014 0.030841 0.1301
9 2002 0.00062 0.055073 0.0682
10 1001 0.00027 0.098345 0.0268
11 364 0.00012 0.175616 0.0076
12 91 5.3E-05 0.3136 0.0015
13 14 2.3E-05 0.56 0.0002
14 1 1E-05 1 1E-05

Now , P(X>3)=1-P(X\leq 3)

=1-[P(X=3)+P(X=2)+P(X=1)+P(X=0)]

=1-(0.0527+0.0168+0.0033+0.0003)

=0.9269

Therefore , the probability that when 14 adults are randomly selected, more than 3 are in excellent health is 0.9269